# Difference-in-Discontinuities: Estimation, Inference and Validity Tests

### v1

@@ D.1 Bandwidth for Testing if Confounding Effects are Constant
This appendix details into the critical issue of bandwidth selection the Wald test to test Assumption3. As mentioned before, this assumption is vital for the difference-in-discontinuitis method, as the non validity means all estimations would be unvalid. The procedure for testing involves considering the following stacked RDDs regression model:
@@
where $T_{-k}$ is a dummy indicating which period the data belongs to, that is, the RDD of which period it’s running. The null hypothesis for the Wald test is:
@@
One central question arises: how should we determine the optimal bandwidth for Equation 8. In this section, we outline the methodology used for bandwidth selection and present simulation results in two different scenarios: one with identical Data Generating Processes for pre- and post-treatment periods and another with different DGPs, both while considering confounding factors. The first scenario aligns with the framework presented in Model 1, as detailed in Appendix D.2.1 while the second encompasses both models 1 and 3 (Appendix D.2.1).
To select bandwidths, we employ the rdrobust package and estimate bandwidths separately for each RDD. Next, we cross the bandwidths and conduct RDD regressions with each of them, evaluating the bias and root-mean-squared errors (RMSE) of the estimated parameters. The chosen bandwidth are the one that minimizes RMSE, alongside the largest and smallest bandwidths, and the CCT (Calonico et al., 2014) bandwidth for all observations regardless of time dummies.
The various bandwidths selected for produced similar results. This finding suggests that the choice of bandwidth does not substantially affect the accuracy of the test. It is essential to note that while this results says this, the choice of bandwidth for the DiDC estimation of the treatment effect remains a crucial consideration of the method.
@@ D.2 Data Generating Processes
We provide further details on each of the data generating processes (DGPs) employed in our simulation studies. Both GDPs employ the same simulation setup based on model 3 from Calonico et al. (2014), with the key difference being the presence of time-fixed confounding factors at the threshold in one of the settings. We perform 1000 replications for each simulation and for each replication, we use a sample size of $n=500$, with
@@ D.2.1 Model 1: Identical DGPs with time-invariant confounders at the threshold
This model includes a time-invariant confounder, c, to represent pre-existing discontinuities at the threshold that affect the outcome $Y_{i}$. Here, units with $Z_{i}\geq 0$ receive a treatment c and units with $Z_{i}<0$ do not. At the same time, units with $Z_{i}\geq 0$ have a different data generating function than units with $Z_{i}<0$. The data generating functions $\mu_{0}(z)$ and $\mu_{1}(z)$, for periods $t=0$ and $t=1$ respectively are as follows
$\mu_{0}(z)$ =$=\begin{cases}0.48+1.27z-0.5\cdot 7.18z^{2}\vspace{-0.5cm}\\ \quad+0.7\cdot 20.21z^{3}+1.1\cdot 21.54z^{4}+1.5\cdot 7.33z^{5}&\text{if }z<0\\ 0.52+0.84z-0.1\cdot 3z^{2}\vspace{-0.5cm}\\ \quad-0.3\cdot 7.99z^{3}-0.1\cdot 9.01z^{4}+3.56z^{5}+\textbf{c}&\text{if }z\geq 0\end{cases}$
$\mu_{1}(z)$ =$=\begin{cases}0.48+1.27z-0.5\cdot 7.18z^{2}\vspace{-0.5cm}\\ \quad+0.7\cdot 20.21z^{3}+1.1\cdot 21.54z^{4}+1.5\cdot 7.33z^{5}&\text{if }z<0\\ 0.52+0.84z-0.1\cdot 3z^{2}\vspace{-0.5cm}\\ \quad-0.3\cdot 7.99z^{3}-0.1\cdot 9.01z^{4}+3.56z^{5}+\textbf{c}+\tau&\text{if}z\geq 0\end{cases}$
@@ D.2.2 Model 2: Identical DGPs
This model follows precisely model 3 from Calonico et al. (2014) for both periods, the only difference between them being that in period $t=1$ there is the effect of the treatment $\tau$ for units with $Z-i\geq 0$. No other discontinuities that could affect the outcome exist at the threshold $z_{0}=0$. The data generating functions $\mu_{0}(z)$ and $\mu_{1}(z)$ are as follows
$\mu_{0}(z)$ =$=\begin{cases}0.48+1.27z-0.5\cdot 7.18z^{2}\vspace{-0.5cm}\\ \quad+0.7\cdot 20.21z^{3}+1.1\cdot 21.54z^{4}+1.5\cdot 7.33z^{5}&\text{if }z<0\\ 0.52+0.84z-0.1\cdot 3z^{2}\vspace{-0.5cm}\\ \quad-0.3\cdot 7.99z^{3}-0.1\cdot 9.01z^{4}+3.56z^{5}&\text{if }z\geq 0\end{cases}$
$\mu_{1}(z)$ =$=\begin{cases}0.48+1.27z-0.5\cdot 7.18z^{2}\vspace{-0.5cm}\\ \quad+0.7\cdot 20.21z^{3}+1.1\cdot 21.54z^{4}+1.5\cdot 7.33z^{5}&\text{if }z<0\\ 0.52+0.84z-0.1\cdot 3z^{2}\vspace{-0.5cm}\\ \quad-0.3\cdot 7.99z^{3}-0.1\cdot 9.01z^{4}+3.56z^{5}+\tau&\text{if }z\geq 0\end{cases}$

### v2

@@ Appendix C Proofs of main results
Proof of Theorem 1.
The result follows from Lemma B.3 by considering estimators from both sides of the threshold together and the conditional independence of observations on either side.
Proof of Theorem 2.
Let $\theta_{0}=\left(\Delta\mu_{+},\Delta\mu_{-}\right)$ and $\widehat{\theta}=\left(\widehat{\mu}_{+,p,q}^{(0)}(h_{n},b_{n}),\widehat{\mu}_{-,p,q}^{(0)}(h_{n},b_{n})\right)$. For any fixed $y$, $c_{1}$ and $c_{2}$, define the mapping
$\phi:l^{\infty}\left(\mathbb{R}\times\mathcal{S}(Z)\right)\times l^{\infty}\left(\mathcal{S}(Z)\right)\rightarrow l^{\infty}\left(\mathcal{S}(Z),\mathbb{R}^{2}\right)$
by
$\left[\phi(\theta)\right](z)=\begin{pmatrix}\max\left\{\theta^{(1)}(y,z)-c_{1},(\theta^{(1)}(y,z)-\theta^{(2)}(y,z))-c_{2}\right\}\\ \min\left\{\theta^{(1)}(y,z)+c_{1},(\theta^{(1)}(y,z)-\theta^{(2)}(y,z))+c_{2}\right\}\end{pmatrix}$
where $\theta^{(j)}$ is the j-th component of $\theta$. From Fang and Santos (2015), the maps $\left(a_{1},a_{2}\right)\mapsto\min\left\{a_{1},a_{2}\right\}$ and $\left(a_{1},a_{2}\right)\mapsto\max\left\{a_{1},a_{2}\right\}$ are Hadamard directionally differentiable with Hadamard directional derivatives at $(a_{1},a_{2})$ equal to
$h\mapsto\left\{\begin{matrix}h^{(1)},if\ a_{1}<a_{2},\\ \min\left\{h^{(1)},h^{(2)}\right\},if\ a_{1}=a_{2},\\ h^{(2)},if\ a_{1}>a_{2}\end{matrix}\right$.
and
$h\mapsto\left\{\begin{matrix}h^{(2)},if\ a_{1}<a_{2},\\ \max\left\{h^{(1)},h^{(2)}\right\},if\ a_{1}=a_{2},\\ h^{(1)},if\ a_{1}>a_{2}\end{matrix}\right$.
for $h\in\mathbb{R}^{2}$.
The mapping comprises max and min operators, along with four Hadamard differentiable functions. We compute their Hadamard derivatives with respect to $\theta$ below:
$\left[\delta_{1}(\theta)\right](z)=\theta^{(1)}(y,z)-c_{1}$ has Hadamard derivative equal to
$\left[\delta^{{}^{\prime}}_{1,\theta}(h)\right](z)=h^{(1)}(y,z)$.
$\left[\delta_{2}(\theta)\right](z)=(\theta^{(1)}(y,z)-\theta^{(2)}(y,z))-c_{2}$ has Hadamard derivative equal to
$\left[\delta^{{}^{\prime}}_{2,\theta}(h)\right](z)=h^{(1)}(y,z)-h^{(2)}(y,z)$.
$\left[\delta_{3}(\theta)\right](z)=\theta^{(1)}(y,z)+c_{1}$ has Hadamard derivative equal to
$\left[\delta^{{}^{\prime}}_{3,\theta}(h)\right](z)=h^{(1)}(y,z)$.
$\left[\delta_{4}(\theta)\right](z)=(\theta^{(1)}(y,z)-\theta^{(2)}(y,z))+c_{2}$ has Hadamard derivative equal to
$\left[\delta^{{}^{\prime}}_{4,\theta}(h)\right](z)=h^{(1)}(y,z)-h^{(2)}(y,z)$.
Using this notation, we write the functional $\phi$ as
$\phi(\theta)=\begin{pmatrix}\max\left\{\delta_{1}(\theta),\delta_{2}(\theta)\right\}\\ \min\left\{\delta_{3}(\theta),\delta_{4}(\theta)\right\}\end{pmatrix}$
By the Chain rule for Hadamard differentiable functions (Lemma C2 from Masten and Poirier, 2020), $\phi$ has Hadamard directional derivative equal to
$\phi_{\theta}^{{}^{\prime}}(h)=\begin{pmatrix}\mathbf{1}\left\{\delta_{1}(\theta)>\delta_{2}(\theta)\right\}\delta_{1,\theta}^{{}^{\prime}}(h)\\ +\mathbf{1}\left\{\delta_{1}(\theta)=\delta_{2}(\theta)\right\}\max\left\{\delta_{1,\theta}^{{}^{\prime}}(h),\delta_{2,\theta}^{{}^{\prime}}(h)\right\}\\ +\mathbf{1}\left\{\delta_{1}(\theta)<\delta_{2}(\theta)\right\}\delta_{2,\theta}^{{}^{\prime}}(h)\\ \\ \mathbf{1}\left\{\delta_{3}(\theta)<\delta_{4}(\theta)\right\}\delta_{4,\theta}^{{}^{\prime}}(h)\\ +\mathbf{1}\left\{\delta_{3}(\theta)=\delta_{4}(\theta)\right\}\min\left\{\delta_{3,\theta}^{{}^{\prime}}(h),\delta_{4,\theta}^{{}^{\prime}}(h)\right\}\\ +\mathbf{1}\left\{\delta_{3}(\theta)>\delta_{4}(\theta)\right\}\delta_{3,\theta}^{{}^{\prime}}(h)\end{pmatrix}$
By theorem B.3, we have $\sqrt{n\min\left\{h_{n},b_{n}\right\}}(\widehat{\theta}(h_{n},b_{n})-\theta_{0})\rightarrow\textbf{Z}(h_{n},b_{n})$, where $\textbf{Z}(h_{n},b_{n})$ is mean-zero gaussian process with variance given by theorem B.3 as well.
Hence, we can use the delta method for Hadamard directionally differentiable functions Fang and Santos (2018) to find that
$\sqrt{n\min\left\{h_{n},b_{n}\right\}}\left[(\phi(\widehat{\theta}(h_{n},b_{n})-\phi(\theta_{0}))\right](z)\rightarrow\left[\phi_{\theta_{0}}^{{}^{\prime}}(\textbf{Z)}\right](z)$
which concludes the proof.
Proof of Lemma 6.
We begin by deriving the sharp identified set for $\tau_{c}$. From Assumption 7, we have
$Y_{i,1}(1,0)+Y_{i,1}(0,1)\leq Y_{i,1}(1,1)+Y_{i,1}(0,0)$
which can be rearranged into
$Y_{i,1}(0,1)-Y_{i,1}(0,0)\leq Y_{i,1}(1,1)-Y_{i,1}(1,0)$
Taking expectations conditional on $Z_{i}=z_{0}$ yields
$\mathbb{E}\left[Y_{i,1}(0,1)-Y_{i,1}(0,0)|Z_{i}=z_{0}\right]\leq\mathbb{E}\left[Y_{i,1}(1,1)-Y_{i,1}(1,0)|Z_{i}=z_{0}\right]=\tau_{c}$
Under Assumptions 1 and 2, these inequality equals
$\mathbb{E}\left[Y_{i,1}(0,1)|Z_{i}=z_{0}\right]-Y_{1}^{-}\leq\tau_{c}$
From assumption 6, it follows that
$y^{min}-Y_{1}^{-}\leq\tau_{c}$
which yields the lower bound for $\tau_{c}$. Note, however, that Assumption 7 provides no improvement on the worst-case upper bound.
We now turn to $\tau_{uc}$. Rearranging Assumption 7 and taking expectations conditional on $Z_{i}=z_{0}$ yields
$\tau_{uc}=\mathbb{E}\left[Y_{i,1}(0,1)-Y_{i,1}(0,0)|Z_{i}=z_{0}\right]\leq\mathbb{E}\left[Y_{i,1}(1,1)-Y_{i,1}(1,0)|Z_{i}=z_{0}\right]$
Under Assumptions 1 and 2, the inequality is equal to
$\tau_{uc}\leq Y_{1}^{+}-E\left[Y_{i,1}(1,0)|Z_{i}=z_{0}\right]$
from which it follows, under Assumption 6, that
$\tau_{uc}\leq Y_{1}^{+}-y^{min}$
which yields the upper bound for $\tau_{uc}$. Note, however, that Assumption 7 provides no improvement on the worst-case lower bound.
Proof of Lemma 7.
We begin by deriving the sharp identified set for $\tau_{c}$. From Assumption 7, we have
$Y_{i,1}(1,0)+Y_{i,1}(0,1)\geq Y_{i,1}(1,1)-Y_{i,1}(0,0)$
which can be rearranged into
$Y_{i,1}(0,1)-Y_{i,1}(0,0)\geq Y_{i,1}(1,1)-Y_{i,1}(1,0)$
Taking expectations conditional on $Z_{i}=z_{0}$ yields
$\mathbb{E}\left[Y_{i,1}(0,1)-Y_{i,1}(0,0)|Z_{i}=z_{0}\right]\geq\mathbb{E}\left[Y_{i,1}(1,1)-Y_{i,1}(1,0)|Z_{i}=z_{0}\right]=\tau_{c}$
Under Assumptions 1 and 2, the inequality equals
$\mathbb{E}\left[Y_{i,1}(0,1)|Z_{i}=z_{0}\right]-Y_{1}^{-}\geq\tau_{c}$
From Assumption 6, it follows that.
$y^{max}-Y_{1}^{-}\geq\tau_{c}$
which yields the upper bound for $\tau_{c}$. Note, however, that there is no improvement on the worst-case lower bound.
We now turn to $\tau_{uc}$. Rearranging Assumption 7 and taking expectations conditional on $Z_{i}=z_{0}$ yields
$\tau_{uc}=\mathbb{E}\left[Y_{i,1}(0,1)-Y_{i,1}(0,0)|Z_{i}=z_{0}\right]\geq\mathbb{E}\left[Y_{i,1}(1,1)-Y_{i,1}(1,0)|Z_{i}=z_{0}\right]$
Under Assumptions 1 and 2, the inequality is equal to
$\tau_{uc}\geq Y_{1}^{+}-E\left[Y_{i,1}(1,0)|Z_{i}=z_{0}\right]$
and from Assumption 6, it follows that
$\tau_{uc}\geq Y_{1}^{+}-y^{max}$
which yields the lower bound for $\tau_{uc}$. Note, however, that there is no improvement on the worst-case upper bound.
Proof of Lemma 8.
We begin with $\tau_{c}$. The inequalities characterizing the bounds in Lemmas 5 and 6 remain valid, which yields the result.
For the case of $\tau_{uc}$, the bounds from Lemma 6 remain valid. Assumption 5 combined with 6 and 7, further implies that
$Y_{i,1}(0,1)-Y_{i,1}(0,0)\leq Y_{i,1}(1,1)-(Y_{i,1}(1,0)-c_{1})$
From the inequality above and Assumptions 1-2, it follows that.
$\tau_{uc}\leq\Delta Y^{+}+c_{1}$
which concludes the proof.
Proof of Lemma 9.
The bounds for $\tau_{c}$ are obtained by combining the bounds from Lemmas 5 and 7.
For the case of $\tau_{uc}$, the bounds from Lemma 7 remain valid. Assumption 5, combined with 6 and 8, further implies that
$Y_{i,1}(0,1)-Y_{i,1}(0,0)\geq Y_{i,1}(1,1)-(Y_{i,0}(1,0)+c_{1})$
From the inequality above and Assumptions 1-2, it follows that.
$\tau_{uc}\geq\Delta Y^{+}-c_{1}$
which concludes the proof.
@@ D.1 Bandwidth for Testing if Confounding Effects are Constant
This appendix details the critical issue of bandwidth selection, the Wald test to test Assumption4. As mentioned before, this assumption is vital for the difference-in-discontinuity method, as the non-validity means all estimations would be invalid. The procedure for testing involves considering the following stacked RDs regression model:
@@
where $T_{-k}$ is a dummy indicating which period the data belongs to, that is, the RD of which period it’s running. The null hypothesis for the Wald test is:
@@
One central question arises: how should we determine the optimal bandwidth for Equation 7. In this section, we outline the methodology for bandwidth selection and present simulation results for two scenarios: one with identical Data Generating Processes for the pre- and post-treatment periods and another with different DGPs, both while accounting for confounding factors. The first scenario aligns with the framework presented in Model 1, as detailed in Appendix D.2.1, while the second encompasses both models 1 and 3 (Appendix D.2.1).
To select bandwidths, we employ the rdrobust package and estimate bandwidths separately for each RD. Next, we cross-validate the bandwidths and conduct RD regressions for each, evaluating the bias and root-mean-squared errors (RMSE) of the estimated parameters. The chosen bandwidth is the one that minimizes RMSE, along with the largest and smallest bandwidths, and the CCT (Calonico et al., 2014b) bandwidth for all observations, regardless of time dummies.
The various bandwidths selected produced similar results. This finding suggests that the choice of bandwidth does not substantially affect the accuracy of the test. It is essential to note that while this result says this, the choice of bandwidth for the DiDC estimation of the treatment effect remains a crucial consideration of the method.
@@ D.2 Data Generating Processes
We provide further details on each of the data-generating processes (DGPs) employed in our simulation studies. Both GDPs employ the same simulation setup based on model 3 from Calonico et al. (2014b), with the key difference being the presence of time-fixed confounding factors at the threshold in one of the settings. We perform 1000 replications for each simulation, and for each replication, we use a sample size of $n=1000$, with
@@ D.2.1 Model 1: Identical DGPs with time-invariant confounders at the threshold
This model includes a time-invariant confounder, c, to represent pre-existing discontinuities at the threshold that affect the outcome $Y_{i}$. Here, units with $Z_{i}\geq 0$ receive a treatment c and units with $Z_{i}<0$ do not. At the same time, units with $Z_{i}\geq 0$ have a different data-generating function than units with $Z_{i}<0$. The data generating functions $\mu_{0}(z)$ and $\mu_{1}(z)$, for periods $t=0$ and $t=1$ respectively are as follows
$\mu_{0}(z)$ $=\begin{cases}0.48+1.27z-0.5\cdot 7.18z^{2}\\ +0.7\cdot 20.21z^{3}+1.1\cdot 21.54z^{4}+1.5\cdot 7.33z^{5}&\text{if }z<0\\ 0.52+0.84z-0.1\cdot 3z^{2}\\ -0.3\cdot 7.99z^{3}-0.1\cdot 9.01z^{4}+3.56z^{5}+\textbf{c}&\text{if }z\geq 0\end{cases}$
$\mu_{1}(z)$ $=\begin{cases}0.48+1.27z-0.5\cdot 7.18z^{2}\\ +0.7\cdot 20.21z^{3}+1.1\cdot 21.54z^{4}+1.5\cdot 7.33z^{5}&\text{if }z<0\\ 0.52+0.84z-0.1\cdot 3z^{2}\\ -0.3\cdot 7.99z^{3}-0.1\cdot 9.01z^{4}+3.56z^{5}+\textbf{c}+\tau&\text{if }z\geq 0\end{cases}$
In addition to Table 1, we present results from Monte Carlo simulations in which the variance changes over time.
Table D.1 shows the results for simulations in a DGP in which the standard deviation of the mean potential outcomes is greater in period 1 than in period 0. We set $\sigma_{\varepsilon}=0.1295$ in period 0 and $\sigma_{\varepsilon}=0.259$ in period 1.
Table D.1: Identical Functional Forms with Confounding (More variance in $t=1$)
AV.Bias RMSE Coverage CIL Bandwidths
$h_{n}$ $b_{n}$
RDD (Robust) 1.044 1.044 0 0.331 0.199 0.331
Diff-in-Disc -0.003 0.074 0.941 0.360 0.221 0.358
NP-DiD 0.001 0.029 0.956 0.314 - -
TWFE 0.001 0.019 1 0.403 - -
Note: Simulations based on 10,000 Monte Carlo experiments with a sample size $n=1000$. RDD is the non-bias corrected RD estimator from CCT (2014), NP-DiD is the outcome regression DiD estimator from Sant’Anna and Zhao (2020), DiDC is the estimator proposed in this paper. “Av. Bias”, “RMSE”, “Cover”, “CIL’ stand for the average simulated bias, simulated root mean-squared errors, 95% coverage probability and 95% confidence interval length, respectively. "Bandwidths" $h_{n}$ and $b_{n}$ report the plug-in bandwidths for point and bias estimation, respectively.
In Table D.2, we invert and set $\sigma_{\varepsilon}=0.259$ in period 0 and $\sigma_{\varepsilon}=0.1295$ in period 1. Overall, the differences in variance over time do not affect the qualitative takeaways from the discussion in Section 6.
Table D.2: Identical Functional Forms with Confounding (More variance in $t=0$)
AV.Bias RMSE Coverage CIL Bandwidths
$h_{n}$ $b_{n}$
RDD (Robust) 1.043 1.043 0 0.175 0.169 0.318
Diff-in-Disc -0.004 0.075 0.929 0.360 0.223 0.360
NP-DiD 0.001 0.029 0.950 0.343 - -
TWFE 0.001 0.018 1 0.404 - -
Note: Simulations based on 10,000 Monte Carlo experiments with a sample size $n=1000$. RDD is the non-bias corrected RD estimator from CCT (2014), NP-DiD is the outcome regression DiD estimator from Sant’Anna and Zhao (2020), DiDC is the estimator proposed in this paper. “Av. Bias”, “RMSE”, “Cover”, “CIL’ stand for the average simulated bias, simulated root mean-squared errors, 95% coverage probability and 95% confidence interval length, respectively. "Bandwidths" $h_{n}$ and $b_{n}$ report the plug-in bandwidths for point and bias estimation, respectively.
@@ D.2.2 Model 2: Identical DGPs
This model follows model 3 from Calonico et al. (2014b) exactly for both periods, the only difference being that in period $t=1$ there is an effect of the treatment $\tau$ for units with $Z_{i}\geq 0$. No other discontinuities that could affect the outcome exist at the threshold $z_{0}=0$. The data generating functions $\mu_{0}(z)$ and $\mu_{1}(z)$ are as follows
$\mu_{0}(z)$ $=\begin{cases}0.48+1.27z-0.5\cdot 7.18z^{2}\\ +0.7\cdot 20.21z^{3}+1.1\cdot 21.54z^{4}+1.5\cdot 7.33z^{5}&\text{if }z<0\\ 0.52+0.84z-0.1\cdot 3z^{2}\\ -0.3\cdot 7.99z^{3}-0.1\cdot 9.01z^{4}+3.56z^{5}&\text{if }z\geq 0\end{cases}$
$\mu_{1}(z)$ $=\begin{cases}0.48+1.27z-0.5\cdot 7.18z^{2}\\ +0.7\cdot 20.21z^{3}+1.1\cdot 21.54z^{4}+1.5\cdot 7.33z^{5}&\text{if }z<0\\ 0.52+0.84z-0.1\cdot 3z^{2}\\ -0.3\cdot 7.99z^{3}-0.1\cdot 9.01z^{4}+3.56z^{5}+\tau&\text{if }z\geq 0\end{cases}$
In addition to Table 2, we present the results for simulations under a DGP in which the functional forms are as displayed above, but the variance changes over time, repeating the exercise from Appendix D.2.1.
Table D.3: Identical Functional Forms without Confounding (More variance in $t=1$)
AV.Bias RMSE Coverage CIL Bandwidths
$h_{n}$ $b_{n}$
RDD (Robust) 0.044 0.079 0.922 0.331 0.199 0.344
Diff-in-Disc -0.003 0.074 0.941 0.360 0.221 0.358
NP-DiD 0.001 0.029 0.956 0.343 - -
TWFE 0.001 0.019 1 0.403 - -
Note: Simulations based on 10,000 Monte Carlo experiments with a sample size $n=1000$. RDD is the non-bias corrected RD estimator from CCT (2014), NP-DiD is the outcome regression DiD estimator from Sant’Anna and Zhao (2020), DiDC is the estimator proposed in this paper. “Av. Bias”, “RMSE”, “Cover”, “CIL’ stand for the average simulated bias, simulated root mean-squared errors, 95% coverage probability and 95% confidence interval length, respectively. "Bandwidths" $h_{n}$ and $b_{n}$ report the plug-in bandwidths for point and bias estimation, respectively.
The results are displayed in Tables D.3 and D.4. Once again, the changes in standard deviations of potential outcomes over time do not affect the qualitative takeaways from Section 6.